Please answer neatly examples....

How can the employment rate rise when the unemployment rate is constant???
Answer #1

Solution:

Such a situation can never occur. Employment and unemployment together make up the labor force. Furthermore,

Denoting Number of people employed by E, no .Of people unemployed by UE, and total number of people in the labor force by LF

Employment rate = E/LF = E/(E + UE)

And unemployment rate = UE/LF = UE/(E + UE)

This means that employment rate + unemployment rate = 1, always.

So, if one of the rates increases, other one has to decrease, or correspondingly, if one of the rates remains same, other has to remain same.

So, if the unemployment rate is constant, situation can never occur that employment rate rises. This is mainly because if the employment rate rises, one of the two simple possible reasons would be that either number of employed has increased highly, while number of unemployed has remained unchanged or decreased slightly (in any of the cases, labor force increases along side) or the number of employed people has remained unchanged, while those unemployed has decreased (in which case labor force decreases). Whatever be the change, it always affects labor force, which further affects the unemployment rate in some way, and hence the unemployment rate will change, which we don't want as per the question.